OpenStudy (anonymous):

find the general solution for dy/dx-y=15 by integration

OpenStudy (inkyvoyd):

dy/dx=y+15

OpenStudy (inkyvoyd):

then you can solve it.

OpenStudy (anonymous):

ln(y+15)=x + c

OpenStudy (inkyvoyd):

@A.Avinash_Goutham , I'm pretty sure that is wrong.

OpenStudy (anonymous):

ok

OpenStudy (inkyvoyd):

@lgbasallote , if you snipe me, I will kill you.

OpenStudy (lalaly):

@A.Avinash_Goutham is right

OpenStudy (inkyvoyd):

NOOOO

OpenStudy (inkyvoyd):

I ragequit OS. I got sniped,

OpenStudy (anonymous):

why not for me it's rigth too

OpenStudy (lgbasallote):

lol inky i didnt say anything

OpenStudy (inkyvoyd):

why isn't it x=y^2+15y+c?

OpenStudy (lalaly):

\[\frac{dy}{y+15}=dx\]integrate both sides\[\ln(y+15)=x+c\]\[e^{\ln(y+15)}=e^{x+c}\]\[y+15=e^{x}e^c\]\[y=Ce^x-15\]

OpenStudy (inkyvoyd):

rage.

OpenStudy (lgbasallote):

differential equations...you cross multiply lol

OpenStudy (anonymous):